# Jr. AI Scientist and Its Risk Report: Autonomous Scientific Exploration from a Baseline Paper

### v1

@@ Abstract
Understanding the current capabilities and risks of AI Scientist systems is essential for ensuring trustworthy and sustainable AI-driven scientific progress while preserving the integrity of the academic ecosystem. To this end, we develop Jr. AI Scientist, a state-of-the-art autonomous AI scientist system that mimics the core research workflow of a novice student researcher: Given the baseline paper from the human mentor, it analyzes its limitations, formulates novel hypotheses for improvement, validates them through rigorous experimentation, and writes a paper with the results. Unlike previous approaches that assume full automation or operate on small-scale code, Jr. AI Scientist follows a well-defined research workflow and leverages modern coding agents to handle complex, multi-file implementations, leading to scientifically valuable contributions. For evaluation, we conducted automated assessments using AI Reviewers, author-led evaluations, and submissions to Agents4Science, a venue dedicated to AI-driven scientific contributions. The findings demonstrate that Jr. AI Scientist generates papers receiving higher review scores than existing fully automated systems. Nevertheless, we identify important limitations from both the author evaluation and the Agents4Science reviews, indicating the potential risks of directly applying current AI Scientist systems and key challenges for future research. Finally, we comprehensively report various risks identified during development. We hope these insights will deepen understanding of current progress and risks in AI Scientist development.
Note: This is a comprehensive report on our AI Scientist project. We do not recommend its use for writing academic papers. Our primary objective is to share insights gained from both successful cases and observed risks with the community to foster a deeper understanding of AI Scientists.
@@ 1 Introduction
As an initial step toward enabling AI Scientists to produce genuine scientific value, we can take inspiration from how student researchers begin their research. When a student first joins a research lab, a common and meaningful process is as follows: the mentor assigns a key paper, the student analyzes its limitations, proposes an improvement hypothesis, implements the idea on the baseline code, validates the hypothesis through thorough experiments, and finally writes a paper summarizing the results. Through this process, the student learns the fundamental workflow of scientific research and gains the skills and experience needed for more creative work later on. Also, improving a baseline method is not only an important stage in early research training but also a valuable research goal in many fields where advancing task performance remains a central scientific pursuit.
@@
As a result of the evaluation using DeepReviewer (zhu2025deepreview), the papers generated by Jr. AI Scientist achieved substantially higher review scores compared to the existing AI-generated papers. Therefore, our Jr. AI Scientist can be regarded as the most capable autonomous AI Scientist. However, we also observed that Jr. AI Scientist still exhibits many failures and unresolved challenges through the author evaluation and Agents4Science conference. To share these challenges and lessons with the research community, we analyze the feedback and evaluation results from the Agents4Science conference and include the author evaluation, which helps clarify what is required to further improve the quality of Jr. AI Scientist systems.
@@
  • Thorough Risk Report: We report the observed risks throughout the project. We believe these reports help deepen the accurate understanding of the current capabilities and limitations of AI Scientist systems.
@@ 2 Related Work
Automated Scientific Discovery. Recent progress has significantly reshaped the role of AI in automating end-to-end scientific research. AI Scientist-v1 (lu2024ai) was an early milestone, showcasing how advanced language models can autonomously generate research ideas, run experiments, and draft scientific papers. This work was followed by a series of subsequent studies (zochi2025, tang2025ai) that further advanced this line of research. However, these approaches often suffer from an overly ambitious problem setting that aims to achieve fully automated science and tend to lack clearly defined scientific objectives for AI Scientists. Without specific goals, such systems often produce undirected discoveries that lack genuine scientific value. To address this issue, our Jr. AI Scientist builds on existing baselines and conducts research within a well-defined research workflow, aiming to generate higher-quality scientific papers. As concurrent work, DeepScientist (weng2025deepscientist) also adopts a baseline-based approach. However, our work differs in that we not only focus on developing state-of-the-art methods, but also aim to contribute to the community by comprehensively reporting the failures and risks encountered throughout the research process.
@@
AI-Assisted Scientific Research. Research specialized for each element of the research process has also been actively explored (chen2025ai4research). For the idea generation phase, si2024can investigates the novelty of the LLM generated ideas, and si2025ideation investigates the ideation–execution gap. For the survey phase, OpenScholar (asai2024openscholar) have been developed to support literature review. For the experimental phase, AlphaEvolve (novikov2025alphaevolve) leverages large-scale trial-and-error strategies to enhance the performance. For the writing and review phase, CycleResearcher (weng2024cycleresearcher) provides a learning framework specialized for scientific writing, while DeepReviewer (zhu2025deepreview) focuses on the review process. Beyond these, rather than pursuing full automation like AI Scientists, AI Co-Scientist (gottweis2025towards) emphasizes collaboration between humans and AI. In this work, instead of focusing on individual parts of the research process, we investigate the entire end-to-end research cycle, aiming to rigorously evaluate both the performance and the limitations.
@@
Failures and Risk Analysis for AI Scientist Systems. There are very few studies that thoroughly analyze or report the risks and failure cases of AI Scientist systems. tang2024prioritizing summarizes the risks that AI Scientists may pose. However, it focuses on hypothesis-based potential risks, rather than empirically observed risks or failures obtained through the actual development and deployment of AI Scientist systems. beel2025evaluating provides an in-depth analysis of failure cases in AI Scientist-v1 (lu2024ai), discussing the low success rate of paper writing, limitations in idea novelty, and the generally poor quality of automatically generated papers. However, these findings are based on the early-stage version of the AI Scientist, and since the analysis was not conducted from the developer’s perspective, the scope of the reported issues is limited. Therefore, we will provide a more comprehensive report on the various risks identified during the development of our state-of-the-art AI Scientist, in order to deepen the community’s understanding of AI Scientists.
@@ 5 Agents4Science Conference Submission
This comment is valid. While Jr. AI Scientist achieved higher scores than the baseline, the performance gap is not significant enough to claim a substantial improvement. To address this limitation, it would be necessary to increase the number of experimental trials and explore more innovative search strategies for selecting experimental nodes.
@@ 7.1 Idea Generation
From this perspective, a more extensive validation was conducted in the concurrent work DeepScientist (weng2025deepscientist), which performed a comprehensive large-scale study. They report that, out of approximately 5,000 unique scientific ideas generated, only 21 ultimately led to genuine scientific innovations. Our experiments require less time because our limitation analysis is effective and our goal is modest, aiming to find one successful idea rather than exploring more successful ideas.

### v2

@@ Abstract
Understanding the current capabilities and risks of AI Scientist systems is essential for ensuring trustworthy and sustainable AI-driven scientific progress while preserving the integrity of the academic ecosystem. To this end, we develop Jr. AI Scientist, a state-of-the-art autonomous AI scientist system that mimics the core research workflow of a novice student researcher: Given the baseline paper from the human mentor, it analyzes its limitations, formulates novel hypotheses for improvement, and iteratively conducts experiments until improvements are realized, and writes a paper with the results. Unlike previous approaches that assume full automation or operate on small-scale code, Jr. AI Scientist follows a well-defined research workflow and leverages modern coding agents to handle complex, multi-file implementations, leading to scientifically valuable contributions. Through our experiments, the Jr. AI Scientist successfully generated new research papers that build upon real NeurIPS, IJCV, and ICLR works by proposing and implementing novel methods. For evaluation, we conducted automated assessments using AI Reviewers, author-led evaluations, and submissions to Agents4Science, a venue dedicated to AI-driven scientific contributions. The findings demonstrate that Jr. AI Scientist generates papers receiving higher review scores than existing fully automated systems. Nevertheless, we identify important limitations from both the author evaluation and the Agents4Science reviews, indicating the potential risks of directly applying current AI Scientist systems and key challenges for future research. Finally, we comprehensively report various risks identified during development. We believe this study clarifies the current role and limitations of AI Scientist systems, offering insights into the areas that still require human expertise and the risks that may emerge as these systems evolve.
Note: This is a comprehensive report on our AI Scientist project. We do not recommend the use of the AI Scientist for generating conference submission papers. Our primary objective is to share insights gained from both successful cases and observed risks with the community to foster a deeper understanding of AI Scientists.
@@ 1 Introduction
As an initial step toward enabling AI Scientists to produce genuine scientific value, we can take inspiration from how student researchers begin their research. When a student first joins a research lab, a common and meaningful process is as follows: the mentor assigns a key paper, the student analyzes its limitations, proposes an improvement hypothesis, implements the idea on the baseline code, iteratively conducts experiments until improvements are realized, and finally writes a paper summarizing the results. Through this process, the student learns the fundamental workflow of scientific research and gains the skills and experience needed for more creative work later on. Also, improving a baseline method is not only an important stage in early research training but also a valuable research goal in many fields where advancing task performance remains a central scientific pursuit.
@@
Through our experiments, the Jr. AI Scientist successfully generated new research papers that build upon the above top-tier venues works by proposing and implementing novel algorithms. As for the evaluation using DeepReviewer (zhu2025deepreview), the papers generated by Jr. AI Scientist achieved substantially higher review scores compared to the existing AI-generated papers. Therefore, our Jr. AI Scientist can be regarded as the most capable autonomous AI Scientist. However, we also observed that Jr. AI Scientist still exhibits some failures and unresolved challenges through the author evaluation and Agents4Science conference. To share these challenges and lessons with the research community, we analyze the feedback and evaluation results from the Agents4Science conference and include the author evaluation, which helps clarify what is required to further improve the quality of Jr. AI Scientist systems.
@@
  • Thorough Risk Report: We report the observed risks throughout the project. We believe these reports offer insights into the areas that still require human expertise and the risks that may emerge as these systems evolve.
@@ 2 Related Work
Automated Scientific Discovery. Recent progress has significantly reshaped the role of AI in automating end-to-end scientific research. AI Scientist-v1 (lu2024ai) was an early milestone, showcasing how advanced language models can autonomously generate research ideas, run experiments, and draft scientific papers. This work was followed by a series of subsequent studies in machine learning fields (zochi2025, tang2025ai) and diverse scientific disciplines (villaescusa2025denario, mitchener2025kosmos) that further advanced this line of research. However, these approaches often suffer from an overly ambitious problem setting that aims to achieve fully automated science and tend to lack clearly defined scientific objectives for AI Scientists. Without specific goals, such systems often produce undirected discoveries that lack genuine scientific value. To address this issue, our Jr. AI Scientist builds on existing baselines and conducts research within a well-defined research workflow, aiming to generate higher-quality scientific papers. As concurrent work, DeepScientist (weng2025deepscientist) also adopts a baseline-based approach. DeepScientist (weng2025deepscientist) focuses mainly on experimental performance, formalizing discovery as a Bayesian Optimization problem. However, its overall framework design and workflow integration are outlined at a high level, with limited discussion of implementation details. In contrast, our Jr. AI Scientist explicitly articulates each stage of the research process and further aims to contribute to the community by comprehensively reporting the failures and risks encountered throughout scientific exploration.
@@
AI-Assisted Scientific Research. Research specialized for each element of the research process has also been actively explored (chen2025ai4research). For the idea generation phase, si2024can investigates the novelty of the LLM-generated ideas, and si2025ideation investigates the ideation–execution gap. For the survey phase, OpenScholar (asai2024openscholar) have been developed to support literature review. For the experimental phase, AlphaEvolve (novikov2025alphaevolve) leverages large-scale trial-and-error strategies to enhance the performance. For the writing and review phase, CycleResearcher (weng2024cycleresearcher) provides a learning framework specialized for scientific writing, while DeepReviewer (zhu2025deepreview) focuses on the review process. Recent studies provide a comprehensive overview of automated review, outlining key challenges, proposing a practical review pipeline for real-world implementation, and constructing a large-scale dataset to support automated review research (ZHUANG2025103332, Lin2023ASPR, lin2023moprd). Beyond these, rather than pursuing full automation like AI Scientists, AI Co-Scientist (gottweis2025towards) emphasizes collaboration between humans and AI. In this work, instead of focusing on individual parts of the research process, we investigate the entire end-to-end research cycle, aiming to rigorously evaluate both the performance and the limitations.
@@
Failures and Risk Analysis for AI Scientist Systems. There are very few studies that thoroughly analyze or report the risks and failure cases of AI Scientist systems. Although tang2024prioritizing summarizes the risks that AI Scientists may pose, it focuses on hypothesis-based potential risks, rather than empirically observed risks or failures. While beel2025evaluating provides an in-depth analysis of failure cases in AI Scientist-v1 (lu2024ai), these findings are based on the early AI Scientist, and the analysis was not conducted from the developer’s perspective. While luo2025more examines four failure modes (benchmark selection, data leakage, metric misuse, and post-hoc bias), their analysis is limited to experimental diagnostics and early AI Scientists without modern coding. Thus, their analysis remains somewhat limited in scope.
Therefore, we will provide a more comprehensive report on the various risks identified during the development of our state-of-the-art AI Scientist, in order to deepen the community’s understanding of AI Scientists.
@@ 5 Agents4Science Conference Submission
While the performance improvement is not necessarily limited, it is indeed true that the Jr. AI Scientist still has considerable room for further enhancement. To achieve higher performance, it would be necessary to increase the number of experimental trials and explore more innovative search strategies for selecting experimental nodes.
@@ 7.1 Idea Generation
From this perspective, more extensive validation was conducted in some recent works (liu2025alphago, weng2025deepscientist). For example, the concurrent work DeepScientist (weng2025deepscientist) performed a comprehensive large-scale study. They report that, out of approximately 5,000 unique scientific ideas generated, only 21 ultimately led to genuine scientific innovations. Our experiments require less time because our limitation analysis is effective, and our goal is modest, aiming to find one successful idea rather than exploring more successful ideas.